An effective rate of 29,61% corresponds to a nominal rate, compounded weekly, of

[1] 29,53%.

[2] 29,61%.

[3] 34,35%.

[4] 26,00%.

[5] none of the above.


1
Expert's answer
2022-02-01T16:23:44-0500

Let effective interest rate=i,

nominal interest rate is j,

"i=100((\frac{1+j}{m})^m-1)",

29.61=100((1+r/52)52-1)).

Put r=26,

29.61=100((1+26/52)52-1)),

29.61=29.61.

LHS=RHS.

Thus, the option 4 is right.